# Kwangwook Seo    Donguk Kwon    Dongha Lee

### v2

@@ Abstract
Recent advancements in table-based reasoning have expanded beyond factoid-level QA to address insight-level tasks, where systems should synthesize implicit knowledge in the table to provide explainable analyses. Although effective, existing studies remain confined to scenarios where a single gold table is given alongside the user query, failing to address cases where users seek comprehensive insights from multiple unknown tables. To bridge these gaps, we propose MT-RAIG Bench, design to evaluate systems on Retrieval-Augmented Insight Generation over Mulitple-Tables. Additionally, to tackle the suboptimality of existing automatic evaluation methods in the table domain, we further introduce a fine-grained evaluation framework MT-RAIG Eval, which achieves better alignment with human quality judgments on the generated insights. We conduct extensive experiments and reveal that even frontier LLMs still struggle with complex multi-table reasoning, establishing our MT-RAIG Bench as a challenging testbed for future research [^1] https://github.com/KWONDU/mt-raig.
@@
Kwangwook Seothanks:   Equal contribution    Donguk Kwon    Dongha Leethanks:   Corresponding author Yonsei University - DLI Lab {tommy2130,2020147570,donalee}@yonsei.ac.kr
@@ 1 Introduction
Although MT-RAIG Bench can serve as a promising testbed for insight-level table reasoning, reliably evaluating long-form outputs in table-based tasks remains a longstanding challenge in the field (Zhao et al., 2024). While recent studies (Zhao et al., 2023b; Wang et al., 2024a) employ table-specific metrics that assess the quality of generated outputs beyond the surface-level matching, they still struggle to evaluate the quality of the insight on the MT-RAIG task. Such limitation arises from their reliance on coarse-grained analyses, which fall short in detecting the finer distinctions needed to check both the output’s groundness on multi-tables and completeness on a multi-hop query.
@@ 2.2 Benchmark Construction
We provide detailed information on the construction of the MT-RAIG Bench. Specifically, we describe the process of creating a machine-annotated dataset, building on efforts from previous works (Tang and Yang, 2024; Wei et al., 2024; Ni et al., 2024; Kim et al., 2024b; Yao et al., 2024) while leveraging several notable advantages over fully human-annotated benchmark. These benefits include: (1) scalability, where large-scale expansion is possible with significantly reduced human labor, and (2) consistency in labeling standards, a crucial aspect that can be more challenging to maintain when multiple human annotators are involved. Additionally, we incorporate human quality checks as critical review points for the automatically generated data, balancing the efficiency of automated annotation with the reliability of human annotation to ensure the high-quality of MT-RAIG Bench. We provide all the detailed prompts in Appendix A.1 and benchmark examples in Appendix E.
@@ 2.2.2 Question Annotation
Unlike existing factoid-level questions, our approach aims to annotate questions that seek comprehensive insights across tables. Following previous work (Kim et al., 2024b), we employ GPT-4o mini as an agent annotator, and generate 10 distinct questions for each table set, designed to capture diverse relational aspects among the tables. To strike a robust middle ground between the reliability of human annotation and the versatility of LLMs, we adopt a human-in-the-loop process that faithfully guides the question annotation while preserving the flexibility of LLMs. Specifically, the question annotation process incorporates three key methods:
@@ 4.2 Results
Table 6: A&R, C&R, P&O, and T&P correspond to the types of question in Table 2. Avg.$\hat{T}_{q}$ denotes the average open-domain performance of different question types with retrieved tables, and Avg.$T_{q}$ indicates closed-domain performance using only the ground truth tables. Precision/Recall for Completeness are detailed in Appendix D.
@@ 6 Conclusion
Lastly, a notable limitation lies in our MT-RAIG Eval’s reliance on LLM for assessment. While recent works increasingly adopt LLM-based evaluators in diverse domains due to their scalability and flexibility (Ye et al., 2024; Kim et al., 2024c; Ru et al., 2024; Han et al., 2024; Wang et al., 2024a), it is important to note that their judgments may still be affected by biases or inconsistencies inherent in the backbone model’s training data or architectural design. However, we believe this issue has been largely minimized in our study. Through a meta-evaluation in Section 3, we demonstrate our MT-RAIG Eval’s reliability in aligning with human judgments. Furthermore, additional experiments that replacing the MT-RAIG Eval’s LLM backbone with open-source models (Table 17) confirm the reproducibility of our evaluation framework.
@@
The output of text generation from LLMs may sometimes contain harmful, biased, or offensive content. However, in our research, we assert that this risk is largely minimized. The source tables and source natural language facts used in the construction of our MT-RAIG Bench are derived from SPIDER (Yu et al., 2018) and Open-WikiTable (Kweon et al., 2023), both of which are publicly available datasets licensed under the CC BY-SA 4.0, and these datasets have been annotated by human experts. Additionally, we manually reviewed the generated texts and eliminate any toxic, offensive, or biased language to ensure the quality and fairness of the auto-generated contents. For human evaluation, two graduate students participated as annotators, receiving a pre-guide before taking part in evaluation and validation. We advise the annotators to complete no more than 20 unit tasks per day, and the entire annotation process spanned approximately 30 days.

### v3

@@ Abstract
Recent advancements in table-based reasoning have expanded beyond factoid-level QA to address insight-level tasks, where systems should synthesize implicit knowledge in the table to provide explainable analyses. Although effective, existing studies remain confined to scenarios where a single gold table is given alongside the user query, failing to address cases where users seek comprehensive insights from multiple unknown tables. To bridge these gaps, we propose MT-RAIG Bench, designed to evaluate systems on Retrieval-Augmented Insight Generation over Multiple-Tables. Additionally, to tackle the suboptimality of existing automatic evaluation methods in the table domain, we further introduce a fine-grained evaluation framework MT-RAIG Eval, which achieves better alignment with human quality judgments on the generated insights. We conduct extensive experiments and reveal that even frontier LLMs still struggle with complex multi-table reasoning, establishing our MT-RAIG Bench as a challenging testbed for future research [^1] https://kwondu.github.io/mt-raig.
@@
Kwangwook Seothanks:   Equal contribution    Donguk Kwon    Dongha Leethanks:   Corresponding author Yonsei University - DLI Lab {tommy2130,donguk.kwon,donalee}@yonsei.ac.kr
@@ 1 Introduction
Although MT-RAIG Bench can serve as a promising testbed for insight-level table reasoning, reliably evaluating long-form outputs in table-based tasks remains a longstanding challenge in the field (Zhao et al., 2024). While recent studies (Zhao et al., 2023b; Wang et al., 2024a) employ table-specific metrics that assess the quality of generated outputs beyond the surface-level matching, they still struggle to evaluate the quality of the insight on the MT-RAIG task. Such limitation arise from their reliance on coarse-grained analyses, which fall short in detecting the finer distinctions needed to check both the output’s groundness on multi-tables and completeness on a multi-hop query.
@@ 2.2 Benchmark Construction
We provide detailed information on the construction of the MT-RAIG Bench. Specifically, we describe the process of creating a machine-annotated dataset, building on efforts from previous works (Tang and Yang, 2024; Wei et al., 2024; Ni et al., 2024; Kim et al., 2024c; Yao et al., 2024) while leveraging several notable advantages over fully human-annotated benchmark. These benefits include: (1) scalability, where large-scale expansion is possible with significantly reduced human labor, and (2) consistency in labeling standards, a crucial aspect that can be more challenging to maintain when multiple human annotators are involved. Additionally, we incorporate human quality checks as critical review points for the automatically generated data, balancing the efficiency of automated annotation with the reliability of human annotation to ensure the high-quality of MT-RAIG Bench. We provide all the detailed prompts in Appendix A.1 and benchmark examples in Appendix E.
@@ 2.2.2 Question Annotation
Unlike existing factoid-level questions, our approach aims to annotate questions that seek comprehensive insights across tables. Following previous work (Kim et al., 2024c), we employ GPT-4o mini as an agent annotator, and generate 10 distinct questions for each table set, designed to capture diverse relational aspects among the tables. To strike a robust middle ground between the reliability of human annotation and the versatility of LLMs, we adopt a human-in-the-loop process that faithfully guides the question annotation while preserving the flexibility of LLMs. Specifically, the question annotation process incorporates three key methods:
@@ 4.2 Results
Table 6: A&S, C&R, P&O, and T&P correspond to the types of question in Table 2. Avg.$\hat{T}_{q}$ denotes the average open-domain performance of different question types with retrieved tables, and Avg.$T_{q}$ indicates closed-domain performance using only the ground truth tables. Precision/Recall for Completeness are detailed in Appendix D.
@@ 6 Conclusion
Lastly, a notable limitation lies in our MT-RAIG Eval’s reliance on LLM for assessment. While recent works increasingly adopt LLM-based evaluators in diverse domains due to their scalability and flexibility (Ye et al., 2024; Kim et al., 2024d; Ru et al., 2024; Han et al., 2024; Wang et al., 2024a), it is important to note that their judgments may still be affected by biases or inconsistencies inherent in the backbone model’s training data or architectural design. However, we believe this issue has been largely minimized in our study. Through a meta-evaluation in Section 3, we demonstrate our MT-RAIG Eval’s reliability in aligning with human judgments. Furthermore, additional experiments that replacing the MT-RAIG Eval’s LLM backbone with open-source models (Table 17) confirm the reproducibility of our evaluation framework.
@@
The output of text generation from LLMs may sometimes contain harmful, biased, or offensive content and contain the risk of potential hallucination (Kim et al., 2024b). However, in our research, we assert that this risk is largely minimized. The source tables and source natural language facts used in the construction of our MT-RAIG Bench are derived from SPIDER (Yu et al., 2018) and Open-WikiTable (Kweon et al., 2023), both of which are publicly available datasets licensed under the CC BY-SA 4.0, and these datasets have been annotated by human experts. Additionally, we manually reviewed the generated texts and eliminate any toxic, offensive, or biased language to ensure the quality and fairness of the auto-generated contents. For human evaluation, two graduate students participated as annotators, receiving a pre-guide before taking part in evaluation and validation. We advise the annotators to complete no more than 20 unit tasks per day, and the entire annotation process spanned approximately 30 days.
Acknowledgement
This work was supported by the IITP grants funded by the Korea government (MSIT) (No. RS-2020-II201361; RS-2024-00457882, AI Research Hub Project), and the NRF grant funded by the Korea government (MSIT) (No. RS-2025-00560295).